# CaRtGS: Computational Alignment for Real-Time Gaussian Splatting SLAM

### v2

@@ IV-B Results
LABEL:Fig:Ablation-Study illustrates our ablation studies conducted on monocular setting of the Replica dataset [26]. These studies were designed to rigorously validate our design choices and substantiate their contributions to the overall system performance. The key findings are listed as follows:
@@
[…]
@@
Opacity regularization is instrumental in reducing the model size without compromising the superior rendering quality. As depicted in LABEL:Fig:opacity-reg, our opacity regularization technique effectively reduces the model size by a factor of two when applying a regularization coefficient of $\lambda_{o}=0.001$. This reduction exerts a minimal negative impact on the PSNR performance, thus maintaining the quality of our renderings. Furthermore, by employing a higher regularization coefficient, for instance, $0.01$, we are able to significantly reduce the number of less critical Gaussian primitives.
V Conclusion
In this work, we introduced CaRtGS, a novel framework that integrates computational alignment with Gaussian Splatting SLAM to achieve real-time photorealistic dense rendering. Our key contribution lies in the development of an adaptive computational alignment strategy that optimizes the rendering process by addressing the computational misalignment inherent in GS-SLAM systems. Through fast splat-wise backpropagation, adaptive optimization, and opacity regularization, we significantly enhanced the rendering quality and computational efficiency of the SLAM process. Looking forward, we envision further improvements by integrating advanced machine learning models to predict and prioritize keyframes for optimization. Additionally, exploring the integration of multi-modal sensors could provide more robust photorealistic scene reconstruction.

### v4

@@ IV-B Results
Figure 6 depicts our ablation studies on the monocular Replica dataset [29], rigorously validating our design choices and highlighting their contributions to system performance. Key findings include:
@@
Adaptive optimization strategically allocates computational resources to enhance rendering quality. Integrating splat-wise backpropagation with adaptive optimization has continuously boosted average PSNR from $33.8$ dB to $34.6$ dB. Furthermore, as illustrated in LABEL:Fig:adaptive-opt, this approach equitably distributes computational resources across keyframes, efficiently addressing long-tail optimization challenges.
@@
Opacity regularization is instrumental in reducing the model size without compromising the superior rendering quality. Our opacity regularization technique, as shown in LABEL:Fig:opacity-reg, can halve the model size with a regularization coefficient of $\lambda_{o}=0.001$, with minimal PSNR performance loss. Increasing the coefficient to $0.01$ further reduces less critical Gaussian primitives, which results in a more efficient model at the expense of some rendering quality.
V Limitations and Feature Work
CaRtGS is an adaptive optimization technology that leverages 3D Gaussian models for high-quality rendering and environmental reconstruction in real-time GS-SLAM systems. Despite its potential, several limitations and challenges are identified below, structured into categories for clarity:
  1. 1.
    Dynamic Environment Challenges. CaRtGS assumes static environments, limiting real-world use and causing tracking failures with dynamic objects.
  2. 2.
    Localization Robustness. CaRtGS focuses on improving the rendering quality of GS-SLAM. However, localization accuracy affects rendering quality, especially in some degeneracy scenarios. Therefore, a robustness localization module is essential for GS-SLAM.
  3. 3.
    Geometry Accuracy. Effective geometry mapping is vital in GS-SLAM. As shown in LABEL:Tab:VECtor, the stereo model’s inferior rendering quality stems from the stereo camera’s suboptimal geometry mapping.
Looking forward, we envision further improvements by integrating advanced machine learning models to predict and handle dynamic objects.
VI Conclusion
In this work, we introduced CaRtGS, a novel framework that integrates computational alignment with Gaussian Splatting SLAM to achieve real-time photorealistic dense rendering. Our key contribution lies in the development of an adaptive computational alignment strategy that optimizes the rendering process by addressing the computational misalignment inherent in GS-SLAM systems. Through fast splat-wise backpropagation, adaptive optimization, and opacity regularization, we significantly enhanced the rendering quality and computational efficiency of the SLAM process.